OpenStudy (anonymous):

(Solve for x) x^3+x^2-2x>0

OpenStudy (amistre64):

factor out the initial common 'x' and then see if the quad leftover is factorable

OpenStudy (amistre64):

it looks to be points of interest at x=0; x = -1/2 +- 3/2

OpenStudy (amistre64):

graph them on a number line and multiply their signage

OpenStudy (amistre64):

<..........-2..........0..........1............> - + + + - - + + - - - + ------------------------------ - + - +

OpenStudy (anonymous):

can u write out the factoring?...im confused

OpenStudy (amistre64):

i could, but its pretty basic enough that i believe you can accomplish it :)

OpenStudy (amistre64):

when we factor out a common x, what are we left with?

OpenStudy (anonymous):

idk. x(x^2+x-2)?

OpenStudy (amistre64):

thats correct; now we just factor apart the quadratic there.

OpenStudy (amistre64):

teh last number gives us a clue; it is made by multiplying two numbers together; what numbers can you multiply to get a 2?

OpenStudy (anonymous):

2 and 1

OpenStudy (amistre64):

very good :) and can we combine 2 and 1 together to get a +1? 2 __ 1 = 1 ; what do we need to put there; + or - ?

OpenStudy (amistre64):

the middle term is a +1x which is why we are doing that

OpenStudy (anonymous):

-

OpenStudy (amistre64):

good; so the quad factors like this: x(x+2)(x-1); now we just need to now what makes each part = 0

OpenStudy (amistre64):

x = 0 when x = ?

OpenStudy (anonymous):

1

OpenStudy (amistre64):

i hope thats a joke.... x = 0 when x equals 0

OpenStudy (amistre64):

1 = 0 is a contradiction

OpenStudy (anonymous):

0 lol

OpenStudy (amistre64):

x+2 = 0 when x = ?

OpenStudy (anonymous):

-2

OpenStudy (amistre64):

good; and for the last one: x-1 = 0 when x = ?

OpenStudy (anonymous):

1

OpenStudy (amistre64):

so those are the points on the number line where x = 0; anything greater than that will be ....well, greater than 0 we map them individually

OpenStudy (amistre64):

<......-2........0..........1.........> - + + + for (x+2) - - + + for x = 0 - - - + for (x-1) ---------------------------- - + - + when they are all multiplied together we get these results

OpenStudy (amistre64):

when x is between -2 and 0; we get a value greater than 0 when x is between 1 and infinity; we get a value greater then 0

OpenStudy (anonymous):

so is it -2>x>1?

OpenStudy (amistre64):

almost.... try that again; we have 2 seperate areas -2 < x < 0 < 1 < x < infinity

OpenStudy (amistre64):

(-2,0) U (1,inf) is another way to write it

OpenStudy (anonymous):

oh ok. but whyd you put the infinity in there?

OpenStudy (amistre64):

<......-2.......0.......1.......> -----( x )---( x )

OpenStudy (amistre64):

because you want all the numbers that make it true; and it is true for all the number from 1 up to as far as you wana count..

OpenStudy (amistre64):

infinity just means that there are no restrictions to get inn the way

OpenStudy (anonymous):

ooohh. dang u are good.

OpenStudy (anonymous):

so that means (-2,0) is in union with (1,inf)

OpenStudy (anonymous):

??

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thanks man u rly helped

OpenStudy (amistre64):

the graph of this thing is like this:

OpenStudy (amistre64):

its greater than 0 when its above the x axis :)

OpenStudy (anonymous):

wait so that that x is greater than 0 when above the x axis?

OpenStudy (anonymous):

means*

OpenStudy (amistre64):

yes; the y axis is the results of the equation

OpenStudy (amistre64):

y is greter than 0 above the xaxis; and less than 0 below it

OpenStudy (amistre64):

the graph is just a visual means of getting the same results

OpenStudy (amistre64):

when x=0; we call those roots of the equation; or zeros to indicate the x intercepts